40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 681     Entry time: Thu Mar 18 15:59:46 2010
Author: Frank 
Type: Laser 
Category: stuff happens 
Subject: Amplifier service 

We tried to adjust the pump light in xtal 1, since it seems to be a little too high - no success, so we aligned it (and the mirror set in front of the amplifier) back.

Note: The original pump current for the diodes was 48 A (so, we set it to this value).

The original temperatures of the pump diodes were:

D1: 20 degC

D2: 20 degC (was: 23.8, so we set it back, since then the output power slightly increased)

D3: 24 degC

D4: 25 degC

The observed output power downstream the PBSs (3) and waveplates (3 HWP + 1 QWP) was 29.5 W.

ELOG V3.1.3-